# Searching for quasinormal modes from Binary Black Hole mergers

### v1

@@ 2 Detection statistic
where the scalar product (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ) is defined by
@@
$\mathcal{Q}[x;\tau_{k},f_{k}]:=\ln\Lambda_{I}[x;\hat{\mathcal{A}},\tau_{k},f_{k}]=\frac{1}{2}N^{-1}MN.$ (10)
@@ 3 Monte Carlo simulations
We have performed Monte Carlo simulations by injecting signals consisting of two modes: $[220]$ (fundamental) and $[330]$ with parameters of the GW190521 event to the LIGO Hanford detector (H1) data. We injected signals at random starting times 2 seconds after the GW190521 merger. We injected signals for an array of signal-to-noise ratios (SNRs) ranging from 5 to 20, and we performed 100 injection for each SNR. For the Kerr method and for SNR $=10$ the 1 standard deviations were $\sigma_{M}=28\,M_{\odot}$ and $\sigma_{a}=0.075$ for estimation of mass and spin respectively. For the agnostic method and SNR $=10$ we obtained 1 standard deviations of $\sigma_{f_{1}}=4$ Hz and $\sigma_{f_{2}}=7.5$ Hz for the two frequencies $f_{1}$ and $f_{2}$ respectively and $\sigma_{\tau_{1}}=9$ ms and $\sigma_{\tau_{2}}=11$ ms for the two damping times $\tau_{1}$ and $\tau_{2}$.
We also note that the waveform of quasinormal modes depends only on the spin $a$ which determines the number of cycles of the signal and not on the mass $M$. Thus our pipeline which was primarily designed for analyzing data from currently operating ground based detectors will also be applicable to analyzing BBH ringdowns in the data of future 3rd generation detectors like Einstein Telescope [7] and Cosmic Explorer [8] and data from space detector LISA [9]. The only difference is that much higher SNRs are expected and consequently more modes can simultaneously be resolved. As an example in Figure 2 we have demonstrated that 4 modes can accurately be resolved for 3rd generation detector with SNR $=75$, and 6 modes for LISA detector where BBH ringdowns with SNRs of even 500 can be available.
Figure 2: Left: 4 modes resolved for 3rd generation detector. SNR $=75$ (mass M = $100\,M_{\odot}$). Right: 6 modes resolved for LISA detector. SNR $=500$ (mass M = $10^{6}\,M_{\odot}$). In both cases spin a $=0.85$.
@@ 4 Case study: GW190521 event ringdown
https://gwosc.org/eventapi/html/GWTC-2.1-confident/GW190521/v4/. For the analysis we preprocessed data by passing it through a narrow band filter with passband [50  130]delimited-[]50130[50\,\,130][ 50 130 ] Hz and whitening. As the starting point of the ringdown we chose the time $t_{0}$ at which the strain of the preprocessed data is maximum. We analyzed data from the two LIGO detectors. There is a rule of thumb that after the time 10 M (M is the total mass of the remnant measured in seconds) the perturbative approximation of the ringdown signal as the sum of damped sinusoids is valid [5]. Consequently we carried out our analysis for an array of ringdowns starting at times $t_{M}=n\times M$ after the time $t_{0}$ where $n$ was varied from 0 to 15. We have first carried out the agnostic search for two damped sinusoids in the ringdown data. The results are presented on the left panel of Figure 3. The analysis found a frequency close to the frequency 68.1 Hz of the dominant quadrupole $[220]$ mode but also revealed another frequency around 100 Hz. The nearby mode to this second frequency is $[330]$ with $f_{[330]}=96.6$ Hz. However we need to keep in mind that there maybe unresolved frequencies near the frequency of the fundamental mode. We see that there is a mode $[210]$ with frequency 57.9 Hz close to the fundamental. We have compared our findings with other analysis available in the literature. We found that the following quasimormal mode compositions were claimed for GW190521 ringdown: $([220],[221],[222])$ — LVC [6], $([220],[330])$ — Capano et al. [10], and $([220],[210]),([220],[210],[320])$ — Siegel et al. [11]. These findings overlap with our results.

### v2

@@ 2 Detection statistic
where the scalar product $(\cdot|\cdot)$ is defined by
@@
$\mathcal{Q}[x;\tau_{k},f_{k}]:=\ln\Lambda_{I}[x;\hat{\mathcal{A}},\tau_{k},f_{k}]=\frac{1}{2}N^{T}M^{-1}N.$ (10)
@@ 3 Monte Carlo simulations
We have performed Monte Carlo simulations by injecting signals consisting of two modes: $[220]$ (fundamental) and $[330]$ with parameters of the GW190521 event to the LIGO Hanford detector (H1) data. We injected signals at random starting times 2 seconds after the GW190521 merger. We injected signals for an array of signal-to-noise ratios (SNRs) ranging from 5 to 20, and we performed 100 injection for each SNR. For the Kerr method and for SNR =10=10 the 1 standard deviations were $\sigma_{M}=28\,M_{\odot}$ and $\sigma_{a}=0.075$ for estimation of mass and spin respectively. For the agnostic method and SNR =10=10 we obtained 1 standard deviations of $\sigma_{f_{1}}=4$ Hz and $\sigma_{f_{2}}=7.5$ Hz for the two frequencies $f_{1}$ and $f_{2}$ respectively and $\sigma_{\tau_{1}}=9$ ms and $\sigma_{\tau_{2}}=11$ ms for the two damping times $\tau_{1}$ and $\tau_{2}$.
We also note that the waveform of quasinormal modes depends only on the spin $a$ which determines the number of cycles of the signal and not on the mass $M$. Thus our pipeline which was primarily designed for analyzing data from currently operating ground based detectors will also be applicable to analyzing BBH ringdowns in the data of future 3rd generation detectors like Einstein Telescope [7] and Cosmic Explorer [8] and data from space detector LISA [9]. The only difference is that much higher SNRs are expected and consequently more modes can simultaneously be resolved. As an example in Figure 2 we have demonstrated that 4 modes can accurately be resolved for 3rd generation detector with SNR =75=75, and 6 modes for LISA detector where BBH ringdowns with SNRs of even 500 can be available.
Figure 2: Left: 4 modes resolved for 3rd generation detector. SNR =75=75 (mass M = $100\,M_{\odot}$). Right: 6 modes resolved for LISA detector. SNR =500=500 (mass M = $10^{6}\,M_{\odot}$). In both cases spin a =0.85=0.85.
@@ 4 Case study: GW190521 event ringdown
https://gwosc.org/eventapi/html/GWTC-2.1-confident/GW190521/v4/. For the analysis we preprocessed data by passing it through a narrow band filter with passband [50  130][50\,\,130] Hz and whitening. As the starting point of the ringdown we chose the time $t_{0}$ at which the strain of the preprocessed data is maximum. We analyzed data from the two LIGO detectors. There is a rule of thumb that after the time 10 M (M is the total mass of the remnant measured in seconds) the perturbative approximation of the ringdown signal as the sum of damped sinusoids is valid [5]. Consequently we carried out our analysis for an array of ringdowns starting at times $t_{M}=n\times M$ after the time $t_{0}$ where $n$ was varied from 0 to 15. We have first carried out the agnostic search for two damped sinusoids in the ringdown data. The results are presented on the left panel of Figure 3. The analysis found a frequency close to the frequency 68.1 Hz of the dominant quadrupole $[220]$ mode but also revealed another frequency around 100 Hz. The nearby mode to this second frequency is $[330]$ with $f_{[330]}=96.6$ Hz. However we need to keep in mind that there maybe unresolved frequencies near the frequency of the fundamental mode. We see that there is a mode $[210]$ with frequency 57.9 Hz close to the fundamental. We have compared our findings with other analysis available in the literature. We found that the following quasimormal mode compositions were claimed for GW190521 ringdown: $([220],[221],[222])$ — LVC [6], $([220],[330])$ — Capano et al. [10], and $([220],[210]),([220],[210],[320])$ — Siegel et al. [11]. These findings overlap with our results.